OpenStudy (anonymous):

We roll a pair of dice. If the sum of the dice is 7, you pay me $26. If the sum is not 7, I pay you the number of dollars indicated by the sum of the dice. What is your expected value for the game?

OpenStudy (whpalmer4):

Let's make a table of all of the ways you can roll the various numbers: 2: 11 3: 12 21 4: 13 22 31 5: 14 23 32 41 6: 15 24 33 42 51 7: 16 25 34 43 52 61 8: 26 35 44 53 62 9: 36 45 54 63 10: 46 55 64 11: 56 65 12: 66 There are 6 ways out of 36 different rolls to roll a 7. 6/36 * ($26) = ($4.33) There is 1 way out of 36 different rolls to roll a 2, which pays $2: 1/36*$2 = $0.06 There are 2 ways out of 36 rolls to roll a 3, paying $3: 2/36*$3 = $0.17 etc. Add them all up (don't forget the $4.33 in parentheses is a negative amount) and that should be your EV.